# CryoSplat: Gaussian Splatting for Cryo-EM Homogeneous Reconstruction

### v2

@@ Appendix A Dataset Details
We provide detailed statistics and characteristics of the four cryo-EM datasets used in our experiments:
@@
  • EMPIAR-10028 (Plasmodium falciparum 80S (Pf80S) ribosome) (Wong et al. 2014): 105,247105{,}247 particle images of size $360\times 360$ pixels at a sampling rate of $1.34$ Å/pixel. This is a widely used benchmark with high-contrast images and a static structure.
@@
  • EMPIAR-10049 (RAG1-RAG2 complex) (Ru et al. 2015): 108,544108{,}544 particles of size $192\times 192$ pixels at $1.23$ Å/pixel. This dataset is considered more challenging due to its lower contrast and flexibility in some regions.
@@
  • EMPIAR-10076 (E. coli large ribosomal subunit undergoing (LSU) assembly) (Davis et al. 2016): 131,899131{,}899 particles of size $320\times 320$ pixels at $1.31$ Å/pixel. This dataset contains substantial conformational and compositional heterogeneity, which poses a challenge to homogeneous modeling.
@@
  • EMPIAR-10180 (Pre-catalytic spliceosome) (Plaschka, Lin, and Nagai 2017): 327,490327{,}490 particles of size $320\times 320$ pixels at $1.69$ Å/pixel. It samples a continuum of conformations, particularly involving large-scale motions of the SF3b subcomplex.
Appendix B Fourier Shell Correlation
@@
To evaluate reconstruction quality on real datasets where ground truth volumes are unavailable, we adopt the gold standard Fourier Shell Correlation (FSC) (Van Heel and Schatz 2005), following established protocols. Each dataset is randomly split into two halves, and the reconstruction algorithm is applied independently to each subset. Let the resulting volumes be $\widehat{\mathcal{V}}_{a}(\mathbf{k})$ and $\widehat{\mathcal{V}}_{b}(\mathbf{k})$, representing their Fourier transforms. The FSC is computed as a function of spatial frequency $k=|\mathbf{k}|$ using the following formula:
$\mathrm{FSC}(k)=\frac{\sum\limits_{|\mathbf{k}|=k}\widehat{\mathcal{V}}_{a}(\mathbf{k})\cdot\widehat{\mathcal{V}}_{b}(\mathbf{k})^{*}}{\sqrt{\Big{(}\sum\limits_{|\mathbf{k}|=k}|\widehat{\mathcal{V}}_{a}(\mathbf{k})|^{2}\Big{)}\Big{(}\sum\limits_{|\mathbf{k}|=k}|\widehat{\mathcal{V}}_{b}(\mathbf{k})|^{2}\Big{)}}}.$ (B.1)
@@
Appendix C More Implementation Details
For all experiments, we do not apply any windowing to the observed particle images during preprocessing. The 3D volume is defined over the domain $E,E]^{3}$, and each 2D projection is assumed to lie within $E,E]^{2}$ in the image plane, where $E=0.5$ defines the spatial extent.
Intuition Behind Initialization
These values are fixed but grounded in straightforward statistical intuition. We observe that most particles are concentrated within a spherical region of radius $E/2$. To reflect this prior and accelerate convergence, we initialize the Gaussian means within this region.
@@ Appendix C More Implementation Details
Backprojection $(D+1)^{3}$ 11 $508$ $396$
CryoDRGN (2021a) $(6\cdot\lfloor D/2\rfloor+L+3)\cdot C$ C=1,024C=1{,}024 11 $680$ 1,0081{,}008
$L\cdot C^{2}+2$ $L=3$ 88 2,5602{,}560 4,9064{,}906
CryoSplat (Ours) $11\cdot N$ N=2,048N=2{,}048 11 $344$ $346$
@@
Table S1: Comparison of GPU memory usage across reconstruction methods at $D=192$ and $D=256$. CryoSplat achieves the lowest and most stable memory footprint, while cryoDRGN incurs significantly higher usage due to its deep neural architecture. All measurements were performed on a single NVIDIA RTX 3090, reflecting the memory usage during the reconstruction of a single full density map. Memory for voxelization and FSC-related half-map reconstruction is not included.
Intuition Behind Learning Rate
In the original 3DGS (Kerbl et al. 2023), different learning rates are assigned to different types of Gaussian parameters (means, scales, rotations, opacities). While this works well in novel view synthesis, it introduces instability in cryo-EM reconstruction. Let the full parameter vector be $\boldsymbol{\theta}=[\mu_{x},\mu_{y},\mu_{z},s_{x},s_{y},s_{z},q_{w},q_{x},q_{y},q_{z},A]^{\top}$. In gradient descent optimization, the direction of parameter updates is determined by the gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}$. Unequal learning rates distort this direction by scaling different components unequally, which can lead to divergence. We observe that such practice causes Gaussians to spread uncontrollably in early iterations and finally diverge. To avoid this, we adopt a single unified learning rate across all parameter types, preserving the intended descent direction and ensuring stable convergence.
Appendix D Additional Results
Memory Usage
@@ Appendix D Additional Results
Tab. S1 presents a comparison of GPU memory usage across different reconstruction methods. CryoDRGN (Zhong et al. 2021a) exhibits the highest memory footprint, exceeding $2.5$ GiB at $D=192$ and approaching $5$ GiB at $D=256$, primarily due to its deep neural decoder and a larger batch size of 8. Interestingly, backprojection consumes more memory at $D=192$ than at $D=256$, which may be attributed to implementation-specific factors such as padding overhead or kernel-level optimizations that favor power-of-two dimensions. This anomaly appears method-specific and does not reflect a general trend. In contrast, cryoSplat demonstrates consistently low memory usage across all configurations. Even with as many as 30,00030{,}000 Gaussians, cryoSplat maintains a GPU memory footprint below $380$ MiB, with negligible variation across resolutions. This efficiency underscores the scalability and suitability of cryoSplat for large-scale or memory-constrained cryo-EM reconstruction scenarios.
@@
Figure S1: Qualitative and quantitative comparison of isotropic and anisotropic GMMs (N=30, [^0] N=30{,}000) on four real datasets. FSC curves show that anisotropic Gaussians consistently achieve higher correlations across spatial frequencies, indicating improved reconstruction accuracy and robustness. Volume visualizations further reveal that anisotropic GMMs better recover fine structural details and directionally extended shapes, whereas isotropic Gaussians tend to fragment such regions.
Isotropic vs. Anisotropic
@@
We investigate how the choice between isotropic and anisotropic Gaussians affects reconstruction quality. While cryoSplat adopts anisotropic Gaussians to represent 3D volumes, it remains compatible with the isotropic formulation commonly used in prior work (2021; 2023; 2023; 2024; 2025). As shown in Fig. S1, anisotropic GMMs achieve higher FSC scores across spatial frequencies and produce sharper, more detailed structures. Subjectively, isotropic Gaussians struggle to capture directionally extended structures and are more prone to noise, often resulting in oversmoothing or distortion. This geometric rigidity may impair their ability to fit complex structures during early optimization, contributing to the unstable convergence observed with random initialization in prior methods. These results highlight the improved representational capacity and reconstruction robustness enabled by anisotropic modeling.
@@
Figure S2: Qualitative evaluation of reconstruction performance with different numbers of Gaussians. Increasing the number of Gaussians leads to visibly improved reconstructions, with finer structural details and enhanced sharpness. Red arrows mark representative regions that highlight the qualitative differences for clearer comparison across settings.
Number of Gaussians
@@
We present visual comparisons of reconstruction results using different numbers of Gaussians. As shown in Fig. S2, increasing the number of components yields progressively sharper and more detailed structures. These qualitative observations align with the quantitative improvements in FSC curves reported in the main paper. Red arrows highlight representative regions where the differences in reconstruction quality are especially pronounced, facilitating direct visual comparison across settings.

### v3

@@ Appendix B Dataset details
We provide detailed statistics and characteristics of the cryo-EM datasets used in our experiments:
@@
  • EMPIAR-10028 (Plasmodium falciparum 80S (Pf80S) ribosome) (Wong et al., 2014): 105,247105{,}247 particle images of size $360\times 360$ pixels at a sampling rate of $1.34$ Å/pixel. This is a widely used benchmark with high-contrast images and a static structure.
@@
  • EMPIAR-10049 (RAG1-RAG2 complex) (Ru et al., 2015): 108,544108{,}544 particles of size $192\times 192$ pixels at $1.23$ Å/pixel. This dataset is considered more challenging due to its lower contrast and flexibility in some regions.
@@
  • EMPIAR-10076 (E. coli large ribosomal subunit undergoing (LSU) assembly) (Davis et al., 2016): 131,899131{,}899 particles of size $320\times 320$ pixels at $1.31$ Å/pixel. This dataset contains substantial conformational and compositional heterogeneity, which poses a challenge to homogeneous modeling.
@@
  • EMPIAR-10180 (Pre-catalytic spliceosome) (Plaschka et al., 2017): 327,490327{,}490 particles of size $320\times 320$ pixels at $1.69$ Å/pixel. It samples a continuum of conformations, particularly involving large-scale motions of the SF3b subcomplex.
  • Synthetic 80S ribosome: We construct a synthetic dataset of the 80S ribosome with 100,000100{,}000 particles using Relion (Scheres, 2016a), following the protocol of Levy et al. (2022a). The electron scattering potential is derived in ChimeraX (Pettersen et al., 2021) at a resolution of $6.0$ Å/pixel, based on two atomic models: the small subunit (PDB 3J7A) and the large subunit (PDB 3J79) (Wong et al., 2014). Each particle image is $128\times 128$ pixels with a pixel size of $3.77$ Å/pixel. Orientations are uniformly sampled over $\mathrm{SO}(3)$, and all images are centered without translations. Defocus values for the CTF are randomly drawn from log-normal distributions following Levy et al. (2022a), and zero-mean white Gaussian noise with varying signal-to-noise ratios (SNRs) is added.
Appendix C Fourier shell correlation
@@
To evaluate reconstruction quality on real datasets without ground truth volumes, we adopt the gold standard Fourier Shell Correlation (FSC) (Van Heel & Schatz, 2005), following established protocols. Each dataset is randomly split into two halves, and the reconstruction algorithm is applied independently to each subset. Let the resulting volumes be $\widehat{V}_{a}({\bm{k}})$ and $\widehat{V}_{b}({\bm{k}})$, representing their Fourier transforms. The FSC is computed as a function of frequency $k$ using the following formula:
$\mathrm{FSC}(k)=\frac{\sum\limits_{\|{\bm{k}}\|_{2}=k}\widehat{V}_{a}({\bm{k}})\cdot\widehat{V}_{b}({\bm{k}})^{*}}{\sqrt{\Big(\sum\limits_{\|{\bm{k}}\|_{2}=k}|\widehat{V}_{a}({\bm{k}})|^{2}\Big)\Big(\sum\limits_{\|{\bm{k}}\|_{2}=k}|\widehat{V}_{b}({\bm{k}})|^{2}\Big)}}.$ (21)
@@
Appendix D More implementation details
D.1 Intuition behind initialization
The values used in initialization are fixed but grounded in straightforward statistical intuition. We observe that most particles are concentrated within a spherical region of radius $E/2$, where $E=0.5$ defines the spatial extent as mentioned in Sec. 4.1. To reflect this prior and accelerate convergence, we initialize the Gaussian means within this region.
@@ D.1 Intuition behind initialization
D.2 Intuition behind learning rate
In the original 3DGS (Kerbl et al., 2023), different learning rates are assigned to different types of Gaussian parameters (means, scales, rotations, opacities). While this works well in novel view synthesis, it introduces instability in cryo-EM reconstruction. Let the full parameter vector be $\bm{\theta}=[\mu_{x},\mu_{y},\mu_{z},s_{x},s_{y},s_{z},q_{w},q_{x},q_{y},q_{z},A]^{\top}$. In gradient descent optimization, the direction of parameter updates is determined by the gradient $\nabla_{\bm{\theta}}\mathcal{L}$. Unequal learning rates distort this direction by scaling different components unequally, which can lead to divergence. We observe that such practice causes Gaussians to spread uncontrollably in early iterations and finally diverge. To avoid this, we adopt a single unified learning rate across all parameter types, preserving the intended descent direction and ensuring stable convergence.
D.3 Optimization algorithm
Our optimization algorithm is summarized in Algorithm 1. Unlike Kerbl et al. (2023), which uses gradient magnitude as the criterion for splitting and cloning Gaussians, we observe that gradients are not a reliable indicator for densification in cryo-EM reconstruction. Furthermore, elaborate densification schemes are generally unnecessary, as our method seldom suffers from significant local minima owing to its close consistency with cryo-EM imaging physics. Nevertheless, we retain a simple densification option to balance efficiency and resolution: fewer Gaussians enable faster training, whereas more Gaussians yield higher-resolution reconstructions, as demonstrated in Fig. 4.
Algorithm 1 Optimization and Densification
NN: number of Gaussians
$D$: side length of the observed particle images
   $\bm{\Theta}$$\leftarrow$ InitAttributes(NN) $\triangleright$ Positions, Scales, Quaternions, Amplitudes
   $i\leftarrow$ 0 $\triangleright$ Epoch Count
   while not converged do
    for ($\bm{Y}$}, $\bm{W}$}, $\bm{t}$}, $\widehat{{\bm{H}}}$) in Dataloader() do $\triangleright$ Observed Image, Rotation, Translation, CTF
     $\bm{Y}$$\leftarrow$ FourierShift($\bm{Y}$}, $\bm{t}$}) $\triangleright$ Center Alignment
     $\bm{Q}$$\leftarrow$ Rasterize($\bm{\Theta}$}, $\bm{W}$}, $D$) $\triangleright$ Algorithm 2
     $\bm{X}$$\leftarrow$ ApplyCTF($\bm{Q}$}, $\widehat{{\bm{H}}}$) $\triangleright$ Apply CTF
     $\mathcal{L}\leftarrow$ Loss($\bm{X}$}, $\bm{Y}$}) $\triangleright$ Loss
     $\bm{\Theta}$$\leftarrow$ Adam($\nabla\mathcal{L}$) $\triangleright$ Backprop and Step
    end for
    if IsDoubleGaussians($i$) then $\triangleright$ (Optional) Densification
     for all Gaussian($\bm{\mu}$}, $\bm{s}$}, $\bm{q}$}, AA) in $\bm{\Theta}$} do
      SplitGaussian($\bm{\mu}$}, $\bm{s}$}, $\bm{q}$}, AA)
     end for
    end if
    $i\leftarrow i+1$
   end while
Algorithm 2 CUDA-accelerated Rasterization
$\bm{\Theta}$}: Gaussian parameters
$\bm{W}$}: viewing transformation matrix
$D$: side length of the observed particle images
   function Rasterize($\bm{\Theta}$}, $\bm{W}$}, $D$)
    $\bm{\mu}$}, $\bm{\Sigma}$}, $A\leftarrow$ BuildGaussians($\bm{\Theta}$})
    $\dot{{\bm{\mu}}}$, $\dot{{\bm{\Sigma}}}\leftarrow$ ViewingTransform($\bm{\mu}$}, $\bm{\Sigma}$}, $\bm{W}$}) $\triangleright$ Viewing Transformation
    $\tilde{{\bm{\mu}}}$, $\tilde{{\bm{\Sigma}}}\leftarrow$ Projection($\dot{{\bm{\mu}}}$, $\dot{{\bm{\Sigma}}}$) $\triangleright$ Orthogonal Projection
    $T\leftarrow$ CreateTiles($D$) $\triangleright$ Tile Count
    $\bm{L}$}, $\bm{K}$$\leftarrow$ DuplicateWithKeys($\tilde{{\bm{\mu}}}$, $T$) $\triangleright$ List of Indices and Keys
    SortByKeys($\bm{K}$}, $\bm{L}$})
    $\bm{R}$$\leftarrow$ IdentifyTileRanges($T$, $\bm{K}$})
    ${\bm{Q}}\leftarrow{\bm{0}}$ $\triangleright$ Init Canvas
    for all Tile $\bm{t}$} in $\bm{Q}$} do
     for all Pixel $\bm{p}$} in $\bm{t}$} do
      $\bm{r}$$\leftarrow$ GetTileRange($\bm{R}$}, $\bm{t}$})
      ${\bm{Q}}({\bm{p}})\leftarrow$ SumSplats($\bm{p}$}, $\bm{L}$}, $\bm{r}$}, $\bm{K}$}, $\tilde{{\bm{\mu}}}$, $\tilde{{\bm{\Sigma}}}$, $A$)
     end for
    end for
    return $\bm{Q}$}
   end function
D.4 Details of the rasterizer
The details of the rasterizer are summarized in Algorithm 2. We follow the tile-based rasterization framework of Kerbl et al. (2023), where the output image is divided into $16\times 16$ pixel tiles, and each splat is instantiated in every tile it overlaps. The splat instances are then assigned keys for sorting, after which each tile can be processed efficiently by locating the corresponding ranges in the sorted list. Since pixels are computed in parallel, the runtime is primarily determined by the maximum number of Gaussians within any tile. For more details, we refer the reader to Kerbl et al. (2023).
Appendix E Additional results
E.1 Multimedia results
We provide videos of rotating reconstructed volumes in the supplementary material for all methods and datasets to facilitate visual comparison.
Table 1: GPU memory usage across reconstruction methods at resolutions ($D=192$, $D=256$).
@@ E.1 Multimedia results
Backprojection $(D+1)^{3}$ 1 508 396
CryoDRGN (Zhong et al., 2021a) $(6\cdot\lfloor D/2\rfloor+L+3)\cdot C$ C=1,024C=1{,}024 1 680 1,008
$L\cdot C^{2}+2$ $L=3$ 8 2,560 4,906
CryoSplat (Ours) $11\cdot N$ N=2,048N=2{,}048 1 344 346
@@
E.2 Memory usage
@@
Tab. 1 compares GPU memory usage across different reconstruction methods. CryoDRGN (Zhong et al., 2021a) exhibits the highest memory footprint, exceeding $2.5$ GiB at $D=192$ and approaching $5$ GiB at $D=256$, primarily due to its deep neural decoder and a larger batch size of 8. Interestingly, backprojection consumes more memory at $D=192$ than at $D=256$, which may be attributed to implementation-specific factors such as padding overhead or kernel-level optimizations that favor power-of-two dimensions. This anomaly appears method-specific and does not reflect a general trend. In contrast, CryoSplat demonstrates consistently low memory usage across all configurations. Even with as many as 30,00030{,}000 Gaussians, CryoSplat maintains a memory footprint below $380$ MiB, with negligible variation across resolutions. This efficiency underscores the scalability and suitability of CryoSplat for large-scale or memory-constrained cryo-EM reconstruction scenarios.
@@
Figure 6: Qualitative and quantitative comparison of isotropic and anisotropic GMMs (N=30, [^0] N=30{,}000) on two datasets (EMPIAR-10028 and EMPIAR-10049). FSC curves show that anisotropic Gaussians consistently achieve higher correlations across spatial frequencies, indicating improved reconstruction accuracy. Volume visualizations further reveal that anisotropic GMMs better recover fine structural details and elongated features, whereas isotropic Gaussians tend to fragment such regions.
E.3 Isotropic vs. anisotropic
CryoSplat represents 3D volumes using anisotropic Gaussians while remaining fully compatible with the isotropic formulation widely adopted in prior works (Chen & Ludtke, 2021; Chen et al., 2023a; b; Schwab et al., 2024; Chen, 2025). When the scaling is isotropic, $s_{x}=s_{y}=s_{z}=\sigma$, the anisotropic Gaussian exactly reduces to the standard isotropic form:
$G({\bm{r}}|{\bm{\mu}},\sigma)=\frac{1}{(2\pi)^{\frac{3}{2}}\sigma^{3}}\exp(-\frac{\|{\bm{r}}-{\bm{\mu}}\|_{2}^{2}}{2\sigma^{2}}),$ (22)
allowing direct integration into existing isotropic GMM-based pipelines.
@@
We investigate the impact of isotropic versus anisotropic Gaussians on reconstruction quality. As shown in Fig. 6, anisotropic GMMs achieve higher FSC scores across spatial frequencies and produce sharper, more detailed structures. Subjectively, isotropic Gaussians struggle to capture elongated features and are often captured by noise, which may contribute to the unstable convergence from random initialization reported in previous methods. These results highlight the improved representational capacity and reconstruction robustness enabled by anisotropic modeling.
E.4 Number of Gaussians
@@
Figure 7: Qualitative evaluation of reconstruction performance with different numbers of Gaussians. Increasing the number of Gaussians leads to visibly improved reconstructions, with finer structural details and enhanced sharpness. Red arrows mark representative regions that highlight the qualitative differences for clearer comparison across settings.
@@
We present visual comparisons of reconstruction results using different numbers of Gaussians. As shown in Fig. 7, increasing the number of components yields progressively sharper and more detailed structures. These qualitative observations align with the quantitative improvements in FSC curves reported in Fig. 3. Red arrows highlight representative regions where the differences in reconstruction quality are especially pronounced, facilitating direct visual comparison across settings.
E.5 Signal-to-noise ratio
Figure 8: Reconstruction performance under varying SNRs. (top) Ground-truth (GT) and reconstructed volumes at different SNR levels. (bottom) Example synthetic particle images corresponding to each SNR. (right) FSC curves between GT and reconstructed volumes across SNRs.
We study the effect of SNR levels on cryoSplat with 5,1205{,}120 Gaussians using the synthetic 80S dataset described in Sec. B. Figure 8 shows example synthetic particles, reconstructed volumes, and FSC curves under varying SNRs. FSCs are computed between the ground truth (GT) and reconstructed volumes. Overall, cryoSplat shows strong noise robustness: SNRs above 0 dB have little impact on reconstruction; high resolution is preserved even under severe noise at 15-15 dB, and reconstructions remain satisfactory at 20-20 dB, despite particles being barely visible.